What is the solution to the equation x(x-1)=0?
Answer by nerdybill(7384)

What is the solution to the equation x(x-1)=0?
.
Set each term on the left to zero and solve:
(x)(x-1)=0
.
So:
x = 0
or
x-1 = 0
x = 1
.
solution: x = {0,1}